# LIME: Localized Image Editing via Attention Regularization in Diffusion Models

### v1

@@ 4.2 Edit Application
IP2P [6] 0.112 0.037 0.852 0.743 0.276 0.158 0.060 0.792 0.618 0.273
IP2P [6] 0.063 0.020 0.933 0.899 0.278 0.096 0.035 0.892 0.827 0.275
Table 1: Evaluation on MagicBrush Dataset [52]. Results for single-turn and multi-turn settings are presented for each method and MB stands for models fine-tuned on MagicBrush. The benchmark values for other approaches are sourced from [52], while values for our proposed method are computed following the same protocol. Across both settings, our method surpasses the base models performance of the compared models. The top-performing is highlighted in bold, while the second-best is denoted with underline for each block.
5 Experiments
5.1 Evaluation Datasets and Metrics
Combining diverse datasets and metrics ensures a thorough evaluation of our proposed method. For each dataset, we report the metrics proposed in the corresponding work.
MagicBrush [52].
The test split offers a comprehensive evaluation pipeline with 535 sessions and 1053 turns. Sessions refer to the source images used for iterative editing instructions, and turns denote the individual editing steps within each session. It employs L1 and L2 norms to measure pixel accuracy, CLIP-I, and DINO embeddings for assessing image quality via cosine similarity, and CLIP-T to ensure that the generated images align accurately with local textual descriptions.
InstructPix2Pix [6].
We evaluate our method on InstructPix2Pix test split with 5K image-instruction pairs. Metrics include CLIP image similarity for visual fidelity and CLIP text-image direction similarity to measure adherence to the editing instructions.
PIE-Bench [20].
EditVal [5].
5.2 Implementation Details
Our method adopts InstructPix2Pix [6] as its base model and runs the model for 100 steps for each stage explained in Secs. 4.1 and 4.2. Specifically, during Edit Localization, intermediate representations are extracted between $30$ and $50$ out of $100$ steps, as suggested in LD-ZNet [34]. Moreover, those intermediate features are resized to $256\times 256$. The number of clusters for segmenting is $8$ across all experiments, motivated by an ablation study. Concurrently, we gather features from steps $1$ to $75$ for the cross-attention maps and retain only related tokens. We extract $100$ pixels with the highest probabilities from the attention maps to identify RoI and determine overlapping segments. For Edit Localization, the image scale $s_{I}$ and the text scale $s_{T}$, the parameters are $1.5$ and $7.5$, respectively. During Edit Application, the attention regularization is employed between steps $1$ and $75$, targeting only unrelated tokens. Throughout the editing process, the image scale, $s_{I}$, and the text scale, $s_{T}$, parameters are set to $1.5$ and $3.5$, respectively.
5.3 Qualitative Results
Figure 4 presents qualitative examples for various editing tasks. These tasks include editing large segments, altering textures, editing multiple small segments simultaneously, and adding, replacing, or removing objects. The first column displays the input images, with the corresponding edit instructions below each image. The second column illustrates the results generated by the base models without our proposed method. The third and fourth columns report the RoI identified by our method and the edited output produced by the base models when our regularization method is applied to these RoIs. As shown in Fig. 4, our method effectively implements the edit instructions while preserving the overall scene context. In all presented results, our method surpasses current state-of-the-art models, including their fine-tuned versions on manually annotated datasets, e.g., MagicBrush [52]. Furthermore, as also claimed and reported in HIVE [53], without additional training, IP2P cannot perform a successful edit for (d) in Fig. 4. However, our proposed method achieves the desired edit without any additional training on the base model as shown Fig. 4 - (d).
5.4 Quantitative Results
Metrics Structure Background Preservation CLIP Similarity
@@ 5.4 Quantitative Results
Methods GT Mask Distance×${}_{{}^{\times 10^{3}}}$downarrow PSNR $\uparrow$ LPIPS×${}_{{}^{\times 10^{3}}}$downarrow MSE×${}_{{}^{\times 10^{4}}}$downarrow SSIM×${}_{{}^{\times 10^{2}}}$uparrow Whole $\uparrow$ Edited $\uparrow$
@@
HIVE [53] + LIME 33.76 24.14 103.63 94.01 81.18 23.62 21.21
HIVE [53] w/MB [52] 34.91 20.85 158.12 227.18 76.47 23.90 21.75
HIVE [53] w/MB [52] + LIME 26.98 26.09 68.28 63.70 84.58 23.96 21.36
HIVE [53] w/MB [52] + LIME 25.86 28.43 50.33 43.25 86.67 24.23 21.43
Table 2: Evaluation on PIE-Bench Dataset [20]. Comparison across ten edit types shows our method outperforming base models on text-guided image editing models. The numbers for the first block are taken from the benchmark paper [20]. The top-performing is highlighted in bold, while the second-best is denoted with underline for each block. GT Mask stands for ground-truth masks as regions of interest.
Results on MagicBrush.
Our method outperforms all other methods on both the single- and multi-turn editing tasks on MagicBrush (MB) [52] benchmark, as seen in  Tab. 1. Compared to the base models, our approach provides significant improvements and best results in terms of L1, L2, CLIP-I, and DINO. For the CLIP-T metric, which compares the edited image and caption to the ground truth, our method comes very close to the oracle scores of $0.309$ for multi-turn and $0.307$ for single-turn. This indicates that our edits accurately reflect the ground truth modifications. VQGAN-CLIP [9] achieves the highest in CLIP-T by directly using CLIP [35] for fine-tuning during inference. However, this can excessively alter images, leading to poorer performance in other metrics. Overall, the performance across metrics shows that our approach generates high-quality and localized image edits based on instructions, outperforming prior state-of-the-art methods.
Results on PIE-Bench.
Our quantitative analysis on the PIE-Bench [20] demonstrates the effectiveness of our proposed method. Compared to baseline models like InstructPix2Pix [6] and fine-tuned versions on MagicBrush [52] and HIVE [53], our method achieves significantly better performance on metrics measuring structure and background preservation. This indicates that our approach makes localized edits according to the instructions while avoiding unintended changes to unaffected regions. At the same time, our method obtains comparable results to base models on the CLIP similarity score, showing it applies edits faithfully based on the textual instruction. A comprehensive comparison is presented in  Tab. 2. Overall, the quantitative results validate that our method can enable text-guided image editing by making precise edits solely based on the given instruction without altering unrelated parts.
Results on EditVal.
In evaluation using the EditVal benchmark dataset [5], our method exhibits superior performance across various edit types, particularly excelling in Object Addition (O.A.), Position Replacement (P.R.), and Positional Addition (P.A.), while achieving second-best in Object Replacement (O.R.). In particular, it performs comparable to other methods for edits involving Size (S.) and Alter Parts (A.P.). A comprehensive comparison is presented in  Tab. 3. Overall, the method advances the state-of-the-art by improving the average benchmark results by a margin of $5\%$ over the previous best model.
@@
IP2P + LIME 0.48 0.49 0.21 0.34 0.49 0.28 0.38
Table 3: Evaluation on EditVal Dataset [5]. Comparison across six edit types shows our method outperforming eight state-of-the-art text-guided image editing models. The numbers for other methods are directly taken from the benchmark paper [5] and the same evaluation setup is applied to our method. The top-performing is highlighted in bold, while the second-best is denoted with underline for each block.
Results on InstructPix2Pix.
We evaluate our method utilizing the same setup as InstructPix2Pix, presenting results on a synthetic evaluation dataset [6] as shown in  Fig. 5. Our approach notably improves the base model, IP2P, optimizing the trade-off between the input image and the instruction-based edit. Additionally, while an increase in text scale, $s_{T}$, enhances the CLIP Text-Image Direction Similarity, it adversely impacts CLIP Image Similarity. For both metrics, the higher, the better. The black arrow indicates the selected configuration for the results in this paper.
Figure 5: InstructPix2Pix Test. Trade-off between input image (Y-axis) and edit (X-axis) is showed. T and C denotes $s_{T}$ and # of clusters, respectively. For all experiments, $s_{I}\in[1.0,2.2]$ is fixed. The arrow points to the chosen configuration for our results.
5.5 Ablation Study
Ablation studies analyze the impact of three key components: the RoI finding method, the number of points from attention maps, and the number of clusters. InstructPix2Pix is the base architecture. Evaluation is on the MagicBrush dataset. Each parameter is modified separately, while other parameters are kept fixed to isolate their impact.
RoI finding methods.
The ground truth masks of MagicBrush [52] are not very tight around the edit area, see Supplementary Material for visualizations. For this reason, our method with predicted masks achieves the best performance for the L1, L2, CLIP-I, and DINO metrics while having on-par results with CLIP-T compared to the use of ground truth masks, as shown in Tab. 4. We also compare the segmentation predicted by adapting the state-of-the-art LPM [33] to IP2P by utilizing the official code base [^1] https://github.com/orpatashnik/local-prompt-mixing. Even in this case, our method achieves better results.
Number of points from attention maps.
Using only $25$ points worsens performance, as it cannot capture multiple distinct segments within RoI. However, having more points includes excessive noise, causing more segments to improperly merge and expanding the RoI area. $100$ points provide better RoI, as shown in Tab. 4.
Number of clusters.
A few clusters like $4$ led to large segments and an expanded RoI, preventing localized edits. Increasing the number of clusters, like $16$ or $32$, causes the separation of a single RoI into multiple clusters. As shown in Tab. 4, $8$ achieves the best results.
Edit Application.
Instead of attention regularization, editing can also be performed in noise space [8, 2, 29]. This corresponds to a linear blending of the input image and a reference image derived from the edit text in noise space, according to the RoI. However, alignment between the reference and input images in the edited area is crucial for targeting the RoI effectively. As shown in Tab. 4 - Edit, our method enhances editing precision by employing attention regularization.
IP2P [6] 0.112 0.037 0.852 0.743 0.276
Mask GT 0.063 0.017 0.935 0.902 0.297
LPM [33] 0.072 0.019 0.924 0.886 0.291
Ours 0.058 0.017 0.935 0.906 0.293
# Points $25$ 0.079 0.023 0.917 0.874 0.290
$100$ 0.058 0.017 0.935 0.906 0.293
$225$ 0.065 0.018 0.932 0.901 0.295
$400$ 0.070 0.020 0.925 0.889 0.295
# Clusters $4$ 0.080 0.022 0.923 0.885 0.295
$8$ 0.058 0.017 0.935 0.906 0.293
$16$ 0.062 0.018 0.933 0.903 0.294
$32$ 0.064 0.018 0.932 0.901 0.291
Edit Noise 0.076 0.022 0.914 0.864 0.291
Ours 0.058 0.017 0.935 0.906 0.293
Table 4: Ablation Study. For fair comparison, all parameters are the same for all settings except the ablated parameter. The top-performing is highlighted in bold, while the second-best is denoted with underline for each block.
6 Conclusion
Input
IP2P [6]
Ours
Input
…it …
…zebra …
Color the tie blue.
Make normal-$\langle$ word normal-$\rangle$ drink from a bucket.
Figure 6: Failure Cases & Limitations. Left: Base model entanglement. Right: Feature mixing issue.
Limitations.
Figure 6 shows limitations of our method: (i) shows the limitation due to the pre-trained base model’s capabilities. Our method can focus on the RoI and successfully apply edits but may alter the scene’s style, particularly in color, due to the base model entanglement. However, our proposal significantly improves the edit compared to IP2P. (ii) illustrates how prompt content impacts edit quality. During editing, all tokens except <start of text>, stop words, and padding, affect the RoI, leading to feature mixing.
Supplementary Material
7 Implementation Details
We obtain the results using an NVIDIA A100 40GB GPU machine with 8 cores. For $512\times 512$ images the IP2P-based baselines (e.g., IP2P, IP2P w/MB, HIVE, and HIVE w/MB) take approximately 15 seconds per edit, while for LIME integrated models, it takes $\approx$25 seconds.
7.1 Baselines
Open-Edit [25]:
This GAN-based approach uses a reconstruction loss for pre-training and incorporates a consistency loss during fine-tuning on specific images. Its unique feature is the arithmetic manipulation of word embeddings within a shared space of visual and textual features.
VQGAN-CLIP [9]:
Enhancing the VQGAN [12] framework with CLIP embeddings [35], this method fine-tunes VQGAN using the similarity of CLIP embeddings between the generated image and the target text, leading to optimized image generation.
SDEdit [27]:
Leveraging the capabilities of Stable Diffusion [38], SDEdit introduces a tuning-free approach. It uses stochastic differential equation noise, adding it to the source image and subsequently denoising to approximate the target image, all based on the target caption.
Text2LIVE [4]:
It propose a Vision Transformer [11] for generating edited objects on an additional layer. It incorporates data augmentation and CLIP [35] supervision, ultimately alpha-blending the edited layer with the original to create the target image.
Null Text Inversion [30]:
By optimizing the DDIM [42] trajectory, this method initially inverts the source image. After, it performs image editing during the denoising process guided by cross-attention [17] between text and image.
SINE [54]:
Real images are edited using model-based guidance and patch-based fine-tuning process.
DreamBooth [39]:
It fine-tunes a diffusion model by learning special text tokens and adjusting model parameters on a set of images for editing.
Textual-Inversion [14]:
It fine-tunes a token embedding within the text-encoder space using a set of images.
Imagic [21]:
It edits images through a three-step process: first fine-tuning a token embedding, then fine-tuning the parameters of a text-guided image diffusion model using the fine-tuned token embedding, and finally performing interpolation to generate various edits based on a target prompt.
DiffEdit [8]:
It identifies the region to edit in images by contrasting between a conditional and unconditional diffusion model based on query and reference texts. Then, it reconstructs the edited image by collecting the features from the text-query by combining the features in the noise/latent space, considering the region to edit.
Blended Latent Diffusion [3]:
This method uses a text-to-image Latent Diffusion Model (LDM) to edit the user-defined mask region. It extracts features for the mask region from the edit text, and for the rest of the image, it uses features from the original image in the noise/latent space.
DirectDiffusion [20]:
It inverts the input image into the latent space of Stable Diffusion [38] and then applies Prompt2Prompt [17] to obtain the desired edit without making any changes to the edit diffusion branch.
Diffusion Disentanglement [48]:
It finds the linear combination of the text embeddings of the input caption and the desired edit to be performed. Since it does not fine-tune Stable Diffusion parameters, they claim that the method performs disentangled edits.
InstructPix2Pix (IP2P) [6]:
Starting from the foundation of Stable Diffusion [38], the model is fine-tuned for instruction-based editing tasks. It ensures that the edited image closely follows the given instructions while maintaining the source image without the need for test-time tuning.
InstructPix2Pix w/MagicBrush [52]:
A version of IP2P [6] trained on MagicBrush train set [52]. Since the MagicBrush dataset has more localized edit examples, the fine-tuned version has better results, as seen in Tab. 1.
HIVE [53]:
It extends IP2P [6] by fine-tuning it with an expanded dataset. Further refinement is achieved through fine-tuning with a reward model, which is developed based on human-ranked data.
HIVE w/MagicBrush [52]:
HIVE [53] fine-tuned on MagicBrush train set [52]. Since the MagicBrush dataset has more localized edit examples, the fine-tuned version has better results, as seen in Tab. 1.
8 Additional Experiments
8.1 MagicBrush Annotations
As mentioned in Sec. 5.5, mask annotations for MagicBrush dataset [52] are not very tight around the edit area, as shown in  Fig. 7. Our method directly uses the identified mask during the editing process, therefore, it is important for the masks to be as tight as possible around the edit area to apply localized edits. The loose GT masks of MagicBrush explain why our model achieves worse performance in  Tab. 4 when using GT masks. To highlight this, we evidentiate with red circles in Fig. 7 precise edits when precise masks are provided to LIME. For the first row - (a), the handle of the racket can be preserved if the mask has a precise boundary between the handle and outfit in the occluded area. Moreover, the second row - (b) shows that if the mask in the MagicBrush dataset is used during the edit, the method changes the color of the blanket as well. However, with the precise mask extracted by our method, the edit can distinguish the objects in the area and apply localized edits.
Input Image
GT
+ Edit
RoI
+ Edit
Make her outfit black.
Can the bed be blue?
8.2 Visual Comparison
VQGAN-CLIP.
As shown in Tab. 1, VQGAN-CLIP [9] has better results on the CLIP-T metric. This is expected since it directly fine-tunes the edited images using CLIP embeddings. However, as seen in Fig. 8, the edited images from VQGAN-CLIP fail to preserve the details of the input image. On the other hand, our method successfully performs the desired edit by preserving the structure and fine details of the scene. It results in similar CLIP-T values with ones for the ground truth edited images in the MagicBrush dataset.

### v2

@@ 5.1 Evaluation Datasets and Metrics
IP2P [7] 0.112 0.037 0.852 0.743 0.276 0.158 0.060 0.792 0.618 0.273
IP2P [7] 0.063 0.020 0.933 0.899 0.278 0.096 0.035 0.892 0.827 0.275
(a) Evaluation on MagicBrush [53]. Results for single- and multi-turn settings are presented for each method and MB stands for models fine-tuned on MagicBrush. The other approaches are sourced from [53], while values for our method are computed by following the same protocol. Across both settings, the integration of LIME surpasses the base model performance, e.g., IP2P and IP2P w/MB.
(b) Evaluation on IP2P Test Split. Trade-off between input image and edit is showed. T and C denotes $s_{T}$ and # of clusters, respectively. For all experiments, $s_{I}\in[1.0,2.2]$ is fixed.
Figure 6: Evalutation on MagicBrush [53] and IP2P Test Split [7].
InstructPix2Pix [7] We evaluate our method on IP2P test split with 5K image-instruction pairs on CLIP image similarity for visual fidelity and CLIP text-image direction similarity to measure adherence to the editing instructions.
5.2 Implementation Details
Our method adopts IP2P [7] as its base model and runs the model for 100 steps for each component in Secs. 4.1 and 4.2. Specifically, during Edit Localization, intermediate representations are extracted between $30$ and $50$ out of $100$ steps, as suggested in LD-ZNet [37]. Moreover, those intermediate features are resized to $256\times 256$. The number of clusters for segmenting is $8$ across all experiments, motivated by Sec. 5.5. We gather cross-attention maps from steps $1$ to $75$ and retain only related tokens. We extract $100$ points with the highest probabilities from the attention maps to identify RoI and determine overlapping segments. For Edit Localization, the image scale $s_{I}$ and the text scale $s_{T}$ are set to $1.5$ and $7.5$, respectively. During Edit Application, the attention regularization is employed between steps $1$ and $75$, targeting only unrelated tokens, e.g., SoT, padding, and stop words. Throughout the editing process, the image scale, $s_{I}$, and the text scale, $s_{T}$, parameters are set to $1.5$ and $3.5$, respectively.
5.3 Qualitative Results
Figure 4 presents qualitative examples for various editing tasks: editing large segments, altering textures, editing multiple small segments simultaneously, and adding, replacing, or removing objects. The first column shows the input images, with the edit instructions below each image. The second column illustrates the results generated by the base models without our proposed method. The third and fourth columns report the RoI identified by our method and the edit result by integrating LIME to base models considering the RoI. As shown in Fig. 4, our method effectively implements the edit instructions while preserving the overall scene. In all presented results, our method surpasses current state-of-the-art models, including their fine-tuned versions on manually annotated datasets, e.g., IP2P w/MB [53].
Figure 5 provides further analyses. The integration of LIME successfully handles diverse tasks, including (a) modifying multiple objects, (b) altering the color of object, (c) adding, (d) removing objects, and altering the texture of an animal (e). As depicted in Fig. 5, our approach significantly surpasses existing baselines. Notably, while they tend to alter entire scene rather than the RoI, e.g., shelf or bird, our method targets and modifies the RoI (a), and (e). Furthermore, baseline models often inadvertently affect areas beyond the intended RoI, as seen in cases (b) and (c). In contrast, our method demonstrates precision by confining its operations within the RoI. Particularly in scenario (d), while baseline models such as IP2P struggle to maintain the integrity of the original image or, as in the case of IP2P w/MB, fail to remove objects effectively, our method excels by accurately removing the specified objects, underscoring its superiority in targeted image editing tasks. Additional comparisons and the applicability of LIME on other methods are provided in Supplementary Material.
5.4 Quantitative Results
Results on MagicBrush Our method outperforms all other methods on both the single- and multi-turn editing tasks on MagicBrush (MB) benchmark [53], as seen in  Fig. 6(a). Compared to the base models, our approach provides significant improvements and best results in terms of L1, L2, CLIP-I, and DINO. For the CLIP-T metric, which compares the edited image and caption to the ground truth, our method comes very close to the oracle scores — metric between GT edited image vs. caption — of $0.309$ for multi-turn and $0.307$ for single-turn. This indicates that our edits accurately reflect the ground truth modifications. VQGAN-CLIP [10] achieves the highest in CLIP-T by directly using CLIP [39] for fine-tuning during inference. However, this can excessively alter images, leading to poorer performance in other metrics. Overall, the performance across metrics shows that our approach generates high-quality and localized image edits based on instructions, outperforming other methods.
Results on InstructPix2Pix We evaluate our method using the same procedure as IP2P, presenting results on a synthetic evaluation dataset [7] as shown in  Fig. 6(b). Our approach notably improves the base model, IP2P, optimizing the trade-off between the input image and edit instruction. Additionally, while an increase in text scale, $s_{T}$, enhances the CLIP Text-Image Direction Similarity, it adversely impacts CLIP Image Similarity. For both metrics, the higher, the better. The black arrow shows the default configuration for our study, which is the basis for this paper’s reporting of all metrics with $s_{T}=3.5$ and the number of clusters is 8.
Table 1: User Study.
Options (Q1) (Q2)
IP2P [7] 22% 25%
+ LIME 72% 65%
Neither 6% 10%
IP2P w/MB [53] 21% 25%
+ LIME 74% 63%
Neither 5% 12%
User Study Although quantitative metrics provide a reasonable means of comparing baselines, evaluating the precision and impact of edits is challenging. Therefore, we conduct a user study, asking 53 anonymous individuals on the crowd-sourcing platform Prolific [38] to answer questions about the localization (Q1) and effectiveness (Q2) of edits for a random selection of 54 pairs of images and edit instructions. Each question has three options: (i) baseline, (ii) baseline + LIME and (iii) none of them (Neither). The findings from our survey indicate that our approach outperforms baselines [7, 53] in both metrics, see Tab. 1.
5.5 Ablation Study
We run ablation studies both on alternative designs for both on components (Sec. 4.1 and Sec. 4.2) of LIME, and on the impact of parameters of LIME.
Table 2: Ablation Study. For all experiments, IP2P is the base architecture and the evaluation is on the MagicBrush dataset. Each parameter is modified separately, while other parameters are kept fixed to isolate their impact.
IP2P [7] 0.112 0.037 0.852 0.743 0.276
Mask Type GT 0.063 0.017 0.935 0.902 0.297
LPM [36] 0.072 0.019 0.924 0.886 0.291
Ours 0.058 0.017 0.935 0.906 0.293
Edit Type Noise 0.076 0.022 0.914 0.864 0.291
Ours 0.058 0.017 0.935 0.906 0.293
# of Points $25$ 0.079 0.023 0.917 0.874 0.290
$100$ 0.058 0.017 0.935 0.906 0.293
$225$ 0.065 0.018 0.932 0.901 0.295
$400$ 0.070 0.020 0.925 0.889 0.295
# of Clusters $4$ 0.080 0.022 0.923 0.885 0.295
$8$ 0.058 0.017 0.935 0.906 0.293
$16$ 0.062 0.018 0.933 0.903 0.294
$32$ 0.064 0.018 0.932 0.901 0.291
Mask type We ablate localization design alternative to Sec. 4.1. Specifically, we employ ground-truth masks provided by MagicBrush [53] and the diffusion-based segmentation from LPM approach [36] integrated into the IP2P framework. The results show that our proposed method outperforms the competing approach on the MagicBrush dataset, see Tab. 2, and performs comparably or better to ground-truth. More analysis is provided in Supplementary Material and further comparisons to alternatively utilizing an off-the-shelf segmentor for localization.
Edit type Instead of attention regularization in  Sec. 4.2, editing can also be performed in noise space [9, 3, 32]. This corresponds to blending of the input image and a reference image derived from the edit text in noise space, based on the RoI. However, alignment between the reference and input images in the edited area is crucial for targeting the RoI effectively. As shown in Tab. 2, our method enhances local editing by employing attention regularization.
Number of points We test with different values to understand the impact of the number of points, $N$, for RoI finding. Using only $25$ points worsens performance, as it cannot capture multiple distinct segments within RoI. However, having more points includes excessive noise, causing more segments to improperly merge and expanding the RoI area. $100$ points provide better RoI, see Tab. 2.
Number of clusters To show the robustness of our method in terms of the number of clusters, we perform an ablation study. A low cluster count (e.g., 4) merges into larger segments, hindering precise edits, while a high count (e.g., 16 or 32) fragments the RoI excessively. Our findings, highlighted in Tab. 2, indicate that 8 clusters strike an optimal balance.
6 Conclusion
Input
IP2P
+ LIME
Input
…it …
…zebra …
Color the tie blue.
Make $\langle$ word $\rangle$ drink from a bucket.
Figure 7: Failure Cases & Limitations.
Limitations.
Our method uses IP2P and inherits its entanglement issue discussed in [7], see Fig. 7 (left), although LIME significantly helps with disentanglement.  Fig. 7 (right) illustrates specific limitations introduced by LIME connected to prompt quality. During the editing process, all tokens except SoT, stop words, and padding, affect the RoI, leading to feature mixing. This limitation could be overcome by altering the edit instruction by using LLMs.
Supplementary Material
We further explore LIME’s applicability to other models like Prompt-to-Prompt [18], HIVE [54], and InstructDiffusion [16], showing improved localized editing through quantitative and qualitative results across datasets like MagicBrush [53], PIE-Bench [21], and EdilVal [5]. We provide ablation studies and implementation details and discuss broader impacts like potential misuse risks balanced against benefits like enhanced creative expression. The material underscores LIME’s ability to enable precise localized image edits while preserving surrounding context.
[…]
A Additional Experiments
A.1 Applicability of LIME to other models
The core concepts behind LIME make it broadly applicable to a variety of image editing models, including Prompt-to-Prompt [18], HIVE [54], and InstructDiffusion [16]. Integrating LIME into these methods offers the potential for enhanced performance in the following ways:
Prompt-to-Prompt
is a prompt-based editing method [18]. Unlike instruction-based editing techniques such as IP2P, an input caption and an output caption are needed to execute the desired image edit. Moreover, since there is no condition on the input image, an image inversion step is required before applying the edit. Prompt-to-Prompt does offer localized editing capabilities, with an extension of the Blend option, which mixes the diffusion processes of two images (original and edited). However, as seen in Fig. 8, integrating the edit application of LIME into Prompt-to-Prompt enables even more precise localized edits. This component, see  Sec. 4.2, ensures that changes are confined to the RoI while preserving the shapes and context of surrounding elements. The official code base [^1] https://github.com/google/prompt-to-prompt/ is used for comparison.
Inverted Image
Prompt-to-Prompt w/o Blend
Prompt-to-Prompt w/ Blend
Prompt-to-Prompt + LIME
Figure 8: The integration of LIME into Prompt-to-Prompt [18]. Red circles and arrows emphasize the localization issues of Prompt-to-Prompt model. The following textual information is used: input caption: A photo of fingernails and output caption: A photo of glitter fingernails.
HIVE
is a fine-tuned version of IP2P on an expanded dataset. Further refinement is achieved through fine-tuning with a reward model, which is developed based on human-ranked data. Table 3 shows the results on the MagicBrush dataset. HIVE [54] improves the performance of IP2P (compare the first and third rows), and further improvement is achieved by fine-tuning HIVE on MagicBrush training set, MB with ✓ stands for it in Tab. 3, (compare the third and fifth rows). For both base HIVE and the version fine-tuned on MB, LIME can further significantly improve performance (compare the third and fourth rows and fifth and sixth rows.).
Table 3: HIVE + LIME Evaluation on MagicBrush [53]. The numbers for others are sourced from [53], while values for our method are computed by following the same protocol.The integration of LIME surpasses the base model performance, e.g., HIVE and HIVE w/MB.
Methods Single-turn Multi-turn
IP2P [7] 0.112 0.037 0.852 0.743 0.276 0.158 0.060 0.792 0.618 0.273
IP2P [7] 0.063 0.020 0.933 0.899 0.278 0.096 0.035 0.892 0.827 0.275
@@ A.1 Applicability of LIME to other models
Fig. 9 presents a qualitative comparison before and after integrating LIME into the HIVE model on samples from the MagicBrush dataset. Figure 9-(a) displays the color change of an object in a scene. While HIVE can effectively implement the edit, it inadvertently alters the structure of the vase and the color of another vase in the background, as highlighted by the red circle. However, with our model integrated, HIVE accurately targets the plants for the desired edit without affecting unrelated areas. In Fig. 9-(b), HIVE fails to recognize one of the fingernails and does not correctly apply the intended edit, indicated by the red circle and arrows. Additionally, HIVE alters the background color to blue, which was mentioned in the edit prompt. In contrast, HIVE, with our model integration, precisely applies edits to the region of interest, such as the fingernails, without missing any parts or altering areas outside the region of interest. Lastly, Fig. 9-(c) demonstrates another instance where HIVE performs an entangled edit, changing the skin color of a woman in the scene when the intended edit was to change the outfit color, as shown by the red arrow. The integration of our model enables localized and separate edits on the input image based on the edit instructions, ensuring that only the specified changes are made.
Input Image
HIVE
HIVE + LIME
Change the plants color to blue.
Put blue glitter on fingernails.
Make her outfit black.
Figure 9: The integration of LIME into HIVE [54]. Red circles and arrows emphasize the localization issues of HIVE model. HIVE + LIME enables localized and effective edits.
InstructionDiffusion (IDiff)
is another IP2P-based method [16]. Integrating LIME into IDiff enhances its performance. IDiff achieved scores of $0.085$, $0.03$, $0.90$, $0.83$, and $0.30$ while the integration of LIME into IDiff improves the scores to $0.071$, $0.02$, $0.92$, $0.86$, and $0.30$ for L1, L2, CLIP-I, DINO, and CLIP-T, respectively, on MagicBrush test dataset. The results can also be comparable with Tab. 3.
Input
IDiff
+ LIME
Instruction: Replace the surfboards with flowers.
A.2 MagicBrush Mask Annotations
As mentioned in Sec. 5.5, the mask annotations for the MagicBrush dataset [53] are not very tight around the edit area which might result in worse edit quality when we use them rather than the segmentation extracted by LIME. We show qualitative results highlighting the problem in  Fig. 10. Our method directly uses the identified mask during the editing process, therefore, it is important for the masks to be as tight as possible around the edit area to apply localized edits. The loose GT masks of MagicBrush explain why our model achieves worse performance in  Tab. 2 when using GT masks. We highlight the significance of precise masks with red circles in Fig. 10. When precise masks are provided to LIME, localized edits can be achieved. For the first row - (a), the handle of the racket can be preserved if the mask has a precise boundary between the handle and outfit in the occluded area. Moreover, the second row - (b) shows that if the mask in the MagicBrush dataset is used during the edit, the method changes the color of the blanket as well. However, with the precise mask extracted by our method, the edit can distinguish the objects in the area and apply localized edits.
Table 4: Evaluation on PIE-Bench [21]. Comparison across ten edit types shows the integration of LIME outperforming base models on instruction-based editing models. GT Mask stands for ground-truth regions of interest masks.
Structure Background Preservation CLIP Similarity
@@ A.2 MagicBrush Mask Annotations
Methods Distance×${}_{{}^{\times 10^{3}}}$downarrow PSNR $\uparrow$ LPIPS×${}_{{}^{\times 10^{3}}}$downarrow MSE×${}_{{}^{\times 10^{4}}}$downarrow SSIM×${}_{{}^{\times 10^{2}}}$uparrow Whole $\uparrow$ Edited $\uparrow$
@@
HIVE w/MB [53] 34.91 20.85 158.12 227.18 76.47 23.90 21.75
HIVE w/MB [53] + LIME 26.98 26.09 68.28 63.70 84.58 23.96 21.36
Input Image
GT
+ Edit
RoI
+ Edit
Make her outfit black.
Can the bed be blue?
These results highlight the quality of the edit masks extracted by our method in an entirely self-supervised way and hint at possible further developments where our contribution could be used to refine the annotations of existing datasets or speed up the creation of new ones.
A.3 More Quantitative Results
PIE-Bench [21]
Quantitative analysis on PIE-Bench [21] demonstrates the effectiveness of our proposed method. Compared to baseline models like IP2P [7] and the fine-tuned version on MagicBrush [53] and HIVE [54], our method achieves significantly better performance on metrics measuring structure and background preservation. This indicates that our approach makes localized edits according to the instructions while avoiding unintended changes to unaffected regions. Edited measures CLIP similarity between edit prompt and edited area, and with Background Preservation provides a measure of localized edits. As seen in Tab. 4, our method is better if both metrics are considered together. At the same time, our method obtains comparable results to base models on the CLIP similarity score, showing that edits are faithfully applied based on the textual instruction. A comprehensive comparison is presented in  Tab. 4. Overall, the quantitative results validate that our method can enable text-guided image editing by making precise edits solely based on the given edit instruction without altering unrelated parts.
EditVal [5]
Table 5: Evaluation on EditVal [5]. Comparison across six edit types shows our method outperforming eight state-of-the-art text-guided image editing models. The numbers for other methods are directly taken from the benchmark dataset [5].
@@ A.3 More Quantitative Results
IP2P [7] + LIME 0.48 0.49 0.21 0.34 0.49 0.28 0.38
Our method exhibits superior performance across various edit types in EditVal benchmark [5], particularly excelling in Object Addition (O.A.), Position Replacement (P.R.), and Positional Addition (P.A.), while achieving second-best in Object Replacement (O.R.). It performs on par with other methods for edits involving Size (S.) and Alter Parts (A.P.). LIME advances the state-of-the-art by improving the average benchmark results by a margin of $5\%$ over the previous best model, see  Tab. 5.
A.4 Visual Comparison to state-of-the-art-methods
A.4.1 VQGAN-CLIP
As shown in Fig. 6(a), VQGAN-CLIP [10] has better results on the CLIP-T metric. This is expected since it directly fine-tunes the edited images using CLIP embeddings. However, as seen in Fig. 11, the edited images from VQGAN-CLIP fail to preserve the details of the input image. On the other hand, our method successfully performs the desired edit by preserving the structure and fine details of the scene and results in a similar CLIP-T score as the one of the ground truth edited images in the MagicBrush dataset.